OpenStudy (anonymous):

5. Micronet stock started the year at $4.40 per share. By June, it was up to $6.82 per share. What percent increased? 6. In July, Micronet dropped to $5.56 per share. What was the percent to decrease from June's price? 7. Micronet dropped to $3.60 a share in October. What was the percent of decrease from July's per share price? 8. By the end of the year, Micronet's price rose back to $4.84 per share. What is the percent of increase from its original price?

OpenStudy (anonymous):

5. Micronet stock started the year at $4.40 per share. By June, it was up to $6.82 per share. What percent increased? 6.82/4.40=1.55 so a 55% increase

OpenStudy (anonymous):

increase decrease increase% = --------- * 100% and decrease% = ---------- * 100% original original 6.82 - 4.40 2.42 5) increase% = ----------- * 100% = ----- * 100% = 55% 4.40 4.40

OpenStudy (anonymous):

You can solve the rest using the formulas I have given

OpenStudy (anonymous):

6. In July, Micronet dropped to $5.56 per share. What was the percent to decrease from June's price? 5.56/6.82=.8152492669, so a 1-.815=18.5% decrease.

OpenStudy (turingtest):

I have walked wangkon through problems 5 and 6, and even set him up the formulas for 7 and 8. He keeps posting this question. He does not want to try to do the problem he just wants answers. I suggest we ignore him.